OpenStudy (anonymous):

(-3)^-3 and write the answer using positive exponents

jimthompson5910 (jim_thompson5910):

\[\large (-3)^{-3}\] \[\large \frac{1}{(-3)^{3}}\] \[\large \frac{1}{-27}\] \[\large -\frac{1}{27}\] So \[\large (-3)^{-3}=-\frac{1}{27}\]